# Conformal Lyapunov Optimization: Optimal Resource Allocation under Deterministic Reliability Constraints

### v1

@@ IV Theoretical Guarantees
In this section, we report theoretical guarantees for the proposed CLO protocol. To this end, we first consider the long-term reliability constraint (II-Ha). As in [12], we make the assumption that the thresholds are upper bounded as
$\theta^{k}_{f}\leq M$ (33)
for some value $M\in[0,1]$ for all times $t$. In practice, the upper bound $M$ can be set to 1 if no further information is available on the threshold sequence $\theta^{k}_{f}$. To obtain a tighter upper bound in (34), the upper bound $M$ on the thresholds can be also estimated after the execution of CLO. The proof of the following claim is given in Appendix A following reference [12].
@@
Under Assumptions 1 and 2, as the number of frames, $F$, grows large, the deterministic long-term reliability constraint (II-Ha) is satisfied by CLO for each realization of the stochastic process $\mathbf{\Omega}(t)$. Specifically, the following upper bound is satisfied by the average reliability loss (23) for any number of frames $F$
$\frac{1}{F}\sum_{f=0}^{F-1}\overline{L_{f}^{k}}\leq\underbrace{r^{k}+\frac{M+\gamma^{k}-\theta^{k}_{0}}{\gamma^{k}F}}_{U(M)}.$ (34)
Proposition 1 shows that, in terms of the reliability constraint, it is advantageous to choose a number of slots per frame, $S$, as small as possible, so as to increase the number of frames $F$ for any given total number of slots $T=FS$. It will be observed next that larger values $S$ are, however, beneficial to reduce the average cost.
The analysis of the cost function in (II-H) and of the average stability constraint (II-Ha) requires the following standard statistical assumption.
@@
be the Lyapunov function for the system’s queues, and assume the condition $\mathbb{E}\{G(fS+1)\}\leq\infty$. Under Assumption 3, denoting by $J_{f}^{*}$ the minimum time average cost at the $f$-th frame achievable by any policy that meets constraint (II-Hb), CLO satisifes the following properties:
@@
where $\mu$ is a constant term.
The role of the Lyapunov function as well as the proof of this results are detailed in Section I of the supplemental materials and B respectively. This proposition shows that CLO can attain a close-to-optimal performance in the long-run, while satisfying all the constraints in problem (II-H). In particular, the suboptimality of the solution is bounded by a term of the order $O(1/S)$. Therefore, improving the network cost requires increasing the frame size $S$.
Overall, the results in this section outline a trade-off in the choice of the number of slots per frame, $S$. In fact, a larger value of $S$ helps obtaining lower levels for the cost function (II-H), while smaller values enhance the speed at which the reliability target (23) is attained.
@@ V Simulation Results
(a) FNR (23) as a function of the time slot index for different frame sizes $S$ $(r^{k}=0.15,\eta=0.5)$.
(b) Energy vs. precision trade-off for different frame sizes $S$.
@@
Figure 7: Percentage of decisions taken at different network nodes for increasing values of the precision penalty parameter $\eta$ in (27).
(a) Energy vs. precision trade-off for LO and CLO.
@@
(c) LO outage probabilities for the worst-case $(M=1)$ and the estimated ($M_{\mathrm{post}}\approx 0.55$) upper bounds in (34).
(d) FNR (23) as a function of the time slot index for LO ($\eta=1$, $r^{k}=0.13$)
Figure 8: Comparisons between LO and CLO: (a) Energy vs. precision trade-offs for LO and CLO; (b) Long-term FNR over time for CLO; (c) Average outage probability over time for CLO; (d) Average FNR over time for LO.
In this section we provide numerical results to test the effectiveness of the proposed CLO protocol and to validate the theoretical guarantees claimed in Sec. IV.
@@ V-A Setting
We consider the single-hop and multi-hop network models depicted in Figure 5. The single-hop network in Figure 5(a) comprises $K=3$ EDs connected to a single edge server. In this scenario, each ED acts as an ES running a UNet segmentation network [47] based on a low complexity minimal MobileNetV3 (MNV3) encoder [48], while the centralized server is equipped with a ResNet50 encoder.
In contrast, in the multi-hop architecture shown in Figure 5(b), there are $|\mathcal{U}|=3$ EDs and $|\mathcal{S}|=4$ servers. The edge and cloud servers are equipped with UNet inference models with increasing complexity as we move from the EDs towards the ES, employing MobileNetV3, ResNet18 and ResNet50 encoders. The models at the different nodes along with their computational complexities are reported in Tables I and II. The network implementations are sourced from the PyTorch Image Models repository [49].
@@
The links between nodes are assumed to be wireless, characterized by a Rayleigh distribution with a path loss $\mathrm{PL}=90\hskip 2.0pt\mathrm{dB}$. We set a maximum transmit power $P_{n}^{\mathrm{max}}=3.5\hskip 2.0pt\mathrm{W}$ for all $n\in\mathcal{N}$, and a noise power spectral density $N_{0}=-174\hskip 2.0pt\mathrm{dBm/Hz}$. All the links are characterized by the same transmission bandwidth $B_{n,m}=20$ MHz for all $(n,m)\in\mathcal{E}$. We set a time time slot duration $\Delta T=50$ ms.
@@ V-B Task Description
We focus on a binary image segmentation task, with images and binary object masks obtained from the Cityscapes dataset [50]. This data set consists of 10,000 training and 10,000 test images of urban scenarios. The images are resized to $256\times 256\times 3$ pixels and encoded in a 32-bit format, resulting in an image size of $W^{k}=768$ KB. Since the Cityscapes dataset is originally designed for multi-class semantic segmentation, we reformulate the task as binary segmentation by labeling only car-related pixels as relevant and treating all others as background. The resulting task models a scenario relevant for vehicular applications.
We impose a reliability constraint on the false negative rate (FNR) (5), while we adopt the ratio between the number of pixels falsely identified as a car over the size of the car, i.e.,
$F_{s}(x,\theta)=\min\left(\frac{\absolutevalue{\overline{y}\cap\mathcal{C}(x,\theta)}}{|y|},1\right),$ (37)
as the precision loss.
To approximate the imprecision loss (see Sec. III-D), we consider a set of low-complexity segmentation architectures based on the PSPNet architecture [51]. Each network corresponds to a specific inference model listed in Tables I and II, and is trained using knowledge distillation techniques to replicate the segmentation mask produced by the teacher network, i.e., the networks actually employed at EDs and ESs. From Tables I, II, and III, we observe that the complexity of the precision predictor takes values in the range $2-7\%$ of the complexity of the original segmentation models.
@@
V-C Precision vs.. Reliability Trade-Off
We start by validating the theoretical guarantees presented in Proposition 1 and Proposition 2. To this end, we assess the impact of different frame sizes $S$ on the trade-offs between energy consumption, precision, and reliability.
@@
Figure 6(a) illustrates the FNR as a function of the time slot index for different frame sizes $S$. Each line corresponds to a single realization. The figure validates the theoretical deterministic guarantees offered by CLO, which are satisfied for every realization.
Figure 6(b) shows the trade-off between overall average precision, evaluated as $1-F_{s}(x,\theta)$, and the total transmission energy consumption. The curves are obtained by varying the penalty $\eta$ in (27) and by averaging over the last 800 time slots for a total number of $T=10,000$ time slots, as well as over $10$ different realizations. Increasing the average precision requires offloading computations to deeper network nodes, which increases transmission energy consumption. Figure 7 supports this observation by showing how decision-making rates vary across different node depths for different values of the penalty hyperparameter $\eta$.
Considering both Figure 6(a) and Figure 6(b) it is observed that, in accordance to Propositions 1 and 2, increasing the frame size $S$ allows CLO to attain a higher precision, while causing a slower convergence of the FNR to the target value.
V-D Comparison with LO
We now compare the performance of CLO with the standard LO, which can only ensure compliance with average reliability constraints. Specifically, with LO, the constraint (23) is replaced by the average requirement
@@ V-D Comparison with LO
LO guarantees this long-term average reliability by reformulating it as a queue stability condition associated with virtual queues $\{Z^{k}\}_{k=1}^{K}$ [8] (see Section II of the supplementary materials for further details).
Accordingly, while CLO updates the reliability hyperparameters $\theta^{k}$ at the end of each time frame, LO treats them as variables that are optimized at each time slot $t$. Treating these variables as discrete with domain within the set $\{0.1,0.2,\dots,0.9\}$ yields a mixed-integer problem, whose complexity grows exponentially with the number of the users $K$. To make the problem computational feasible, we force all the users to employ the same threshold, i.e., $\theta^{k}(t)=\theta^{*}(t)$ for all $k$ and $t$ for LO.
@@
The main interest in comparing LO and CLO is to understand the price that CLO must inevitably pay to guarantee a deterministic, per-realization, reliability constraint. To elaborate on this, Figure 8(a) compares the average precision achieved by LO and CLO as a function of the energy consumption. These results are evaluated at convergence of the reliability constraint, averaging over the last 1000 slots for a total number of $T=15,000$ time slots, and considering a single realization. LO is observed to achieve a higher precision for the same energy consumption as compared to CLO, with the gap quantifying the cost accrued by CLO to ensure reliability constraints.
@@
The reliability constraints are highlighted in Figures 8(b) and 8(d), which show the FNR as a function of the time slot index for 10 realizations of LO and CLO. These plots are obtained under comparable energy consumption for both the optimization strategies, which corresponds to the rightmost points in Figure 8(a). The continuous blue curve in Figures 8(b) and 8(d) corresponds to the worst-case upper bound, computed by setting $M=1$ in (34), while the red curve corresponds to an a posteriori upper bound, obtained by estimating the value of the constant $M$ among 10 realizations of CLO.
@@
In Figure 8(c), we present the average outage probabilities for LO, corresponding to the worst case and to the a posteriori estimated upper bound. The curves are obtained by evaluating the fraction of realizations, among a total of 50 realizations, that given an FNR level above the upper bound in Proposition 1, which is instead guaranteed by CLO. While CLO consistently remains within the theoretical bounds, LO exhibits a high likelihood of exceeding them, with this probability increasing over time as the constraint tightens.

### v5

@@ IV Theoretical Guarantees
In this section, we report theoretical guarantees for the proposed CLO protocol. To this end, we first consider the long-term reliability constraint (II-Ha). The proof of the following claim is given in Appendix B following reference [12].
@@
Under Assumptions 1 and 2, as the number of frames, $F$, grows large, the deterministic long-term reliability constraint (II-Ha) is satisfied by CLO for each realization of the stochastic process $\mathbf{\Omega}(t)=\{\mathbf{A}(t),\mathbf{S}(t),\mathbf{T}(t)\}$. Specifically, the following lower and upper bounds $l(m)$, $U(M)$, are satisfied by the average reliability loss (23) for any number $F$ of frames
$l(m)$ $=r^{k}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{r^{k}d^{k}}{F}}+\frac{m-\gamma^{k}-\theta^{k}_{0}}{\gamma^{k}F},$
$U(M)$ $=r^{k}+\frac{M+\gamma^{k}-\theta^{k}_{0}}{\gamma^{k}F}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{d^{k}(1-r^{k})}{F}},$
$\quad l(m)$ $\leq\frac{1}{F}\sum_{f=0}^{F-1}\overline{L_{f}^{k}}\leq U(M).$ (33)
where $M=\max_{f}\{\theta^{k}_{f}\}-\gamma^{k}$ and $m=\min_{f}\{\theta^{k}_{f}\}+\gamma^{k}$.
For classification and binary segmentation tasks, we can set $M=1$ and $m=0$ if no further information is available [12]. Otherwise, the value of $M$ ($m$) can be estimated from the maximum (minimum) hyperparameter $\theta^{k}_{f}$ observed after the execution of CLO, thus obtaining tighter (a posteriori) bounds.
Proposition 1 shows that, in terms of the reliability constraint, it is advantageous to choose a number of slots per frame, $S$, as small as possible, so as to increase the number of frames $F$ for any given total number of slots $T=FS$. However, it will be observed next that larger values of $S$ are beneficial to reduce the average cost.
The analysis of the cost function in (II-H), and of the average stability constraint (II-Ha), requires the following standard statistical assumption.
@@
be the Lyapunov function for the system’s queues, and assume the condition $\mathbb{E}\{G(fS+1)\}\leq\infty$. Under Assumption 3, denoting by $J_{f}^{*}$ the minimum time-average cost at the $f$-th frame achievable by any policy that meets constraint (II-Hb), CLO satisfies the following properties:
@@
where $\mu$ is a constant term, and $V$ is the LO hyperparameter that trades the minimization of the objective function (i.e., energy and precision loss) for the average system delay.
The role of the Lyapunov function, as well as the proof of this results, are detailed in Section I of the supplemental materials and Appendix C, respectively. This proposition shows that CLO can attain a close-to-optimal performance in the long-run, while satisfying all the constraints in problem (II-H). In particular, the sub-optimality of the solution is bounded by a term of the order $\mathcal{O}(1/S)$. Therefore, improving the network cost requires increasing the frame size $S$.
Overall, the results in this section outline a trade-off in the choice of the number $S$ of slots per frame. In fact, a larger value of $S$ helps obtaining lower levels for the cost function (II-H), while smaller values enhance the speed at which the reliability target (23) is attained.
@@ V Simulation Results
(a) FNR (23) as a function of the time slot index for different frame sizes $S$ $(r^{k}=0.15,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d^{k}=0},\eta=0.5)$.
(b) Energy vs. precision trade-off for different frame sizes $S$ ($d^{k}=0$).
@@
Figure 7: Percentage of decisions at different network nodes for different precision-energy trade-off parameter $\eta$ in (26), ($S=50$, $d^{k}=0$).
(a) Energy vs. precision trade-off for CLO, LO, and LO with outage probability constraints.
@@
(c) LO outage probabilities for the worst-case $(M=1)$ and the estimated ($M_{\mathrm{post}}\approx 0.6$) upper bounds in (33).
(d) FNR (23) as a function of the time slot index for LO ($\eta=1$, $r^{k}=0.13$).
Figure 8: Comparisons between LO and CLO assuming no delay in estimation and dissemination of frame-loss information (i.e., $d^{k}=0$): (a) Energy vs. precision trade-offs for LO and CLO; (b) Long-term FNR over time for CLO; (c) Average outage probability over time for CLO; (d) Average FNR over time for LO.
In this section we provide numerical results to test the effectiveness of the proposed CLO protocol and to validate the theoretical guarantees claimed in Section IV.
@@ V-A Setting
We consider both a single-hop and a multi-hop network, as summarized in Figure 5. The single-hop network in Figure 5(a) comprises $K=3$ EDs connected to a single centralized ES, which is equipped with a ResNet50 encoder. Each ED acts also as an ES running a UNet segmentation network [47] based on a minimal (e.g., low complexity) MobileNetV3 (MNV3) encoder [48].
In contrast, in the multi-hop architecture shown in Figure 5(b), there are $|\mathcal{U}|=3$ EDs and $|\mathcal{S}|=4$ servers. The edge and cloud servers are equipped with UNet inference models, characterized by an increasing complexity (and possibly higher precision), as we move from the EDs towards the edge/cloud servers, employing MobileNetV3, ResNet18 and ResNet50 encoders. The NNs employed at each node for image segmentation, along with their computational complexities, are reported in Tables I and II. Their implementation exploits the PyTorch Image Models repository [49].
@@
The links between nodes are assumed to be wireless and, for simplicity, characterized by a Rayleigh distribution with the same average path-loss $\mathrm{PL}=90\hskip 2.0pt\mathrm{dB}$. We set a maximum transmit power $P_{n}^{\mathrm{max}}=3.5\hskip 2.0pt\mathrm{W}$ for all the nodes $n\in\mathcal{N}$, and the same noise power spectral density $N_{0}=-174\hskip 2.0pt\mathrm{dBm/Hz}$. All the links are characterized by the same transmission bandwidth $B_{n,m}=20$ MHz for all $(n,m)\in\mathcal{E}$. We set a time time slot duration $\delta=50$ ms, corresponding to the channel coherence time. We assume that there is no delay associated with the estimation and dissemination of frame loss information, i.e., $d^{k}=0$ for all EDs (see Appendix D for further results). The per-slot problem (31) is solved using the Python-based CVX implementation (CVXPY) [50, 51].
@@ V-B Task Description
We focus on a binary image segmentation task, with images and binary object masks obtained from the Cityscapes dataset of urban scenarios [52]. We split this data set in 10,000 images for training, and 10,000 images for testing the segmentation NNs. The images are resized to $256\times 256\times 3$ pixels and encoded in a 32-bit format, resulting in an image size of $W^{k}=768$ KB. Since the dataset is originally designed for multi-class semantic segmentation, we formulate the task as a binary segmentation by labeling only car-related pixels as segmentation objective, treating all the others as background. For instance, this task could be useful in vehicular applications.
We set $r^{k}=0.15$ in (23) for the FNR constraint, while for the precision loss $F_{s}(x,\theta)$ we consider the ratio of the pixels falsely identified as part of the car, over the true ones  [^3] We chose this precision measure because, unlike (9), it is relative to the size of the object of interest, making it more meaningful for objects that are significantly smaller than the background., i.e.,
$F_{s}(x,\theta)=\min\left(\frac{\absolutevalue{\overline{y}_{\mathrm{true}}}\cap\mathcal{C}(x,\theta)}{|y_{\mathrm{true}}|},1\right).$ (36)
To predict the precision loss $F_{s}(x,\theta)$ (see Section III-D), we consider a set of low-complexity NNs based on the PSPNet architecture [53]. Each NN approximates the precision loss $F_{s}(x,\theta)$ of the inference model employed at the $s$-th ES, which are summarized in Tables I and II. These NNs are trained using knowledge distillation [46], by minimizing a linear combination of the segmentation loss (i.e., cross-entropy) and the Kullback–Leibler divergence between the outputs of the teacher and student NNs. This approach enables the student NNs to replicate the segmentation masks produced by the teacher NNs, i.e., the models actually deployed at the EDs and ESs. From Tables I, II, and III, we observe that the complexity of the precision predictors (PP) takes values in the range $2\%-7\%$ of the complexity of the actual segmentation models.  [^4] Thus the use of the PPs at the (unique) control center makes sense because it requests a much lower complexity than directly performing the segmentation assigned to the ESs.
@@
V-C Precision-Reliability Trade-Off
We start by validating the theoretical guarantees presented in Proposition 1 and Proposition 2, by assessing the impact of a different frame size $S$ on the trade-offs between energy consumption, precision, and reliability.
@@ V-C Precision-Reliability Trade-Off
Figure 6(a) plots the FNR evolution in time for different frame sizes $S$, where each curve is obtained for the same single realization of tasks. The figure shows the theoretical deterministic guarantees offered by CLO.
Figure 6(b) shows the trade-off between overall average precision, evaluated as $1-F_{s}(x,\theta)$, and the transmission energy consumption of all the nodes. The curves are obtained by varying the penalty $\eta$ in (26) and by averaging over the last 1,000 time slots (of the total $T=10,000$), as well as over $30$ different realizations of tasks. Increasing the average precision requires offloading computations to network nodes farther away from the users, increasing the transmission energy consumption. This is confirmed by Figure 7, which shows how the nodes decision percentages vary with the nodes depth, for different values of the precision-energy trade-off parameter $\eta$.
Figures 6(a) and 6(b) confirm that, in accordance to Propositions 1 and 2, increasing the frame size $S$ allows CLO to attain a higher precision over a finite time duration, but a slower convergence to the FNR target value.
V-D Comparison with LO-based Resource Allocation Strategies
Figure 9: Long-term FNR reliability loss for CLO, LO, and LO with outage probability constrains $(\eta=1,r^{k}=0.13,d^{k}=0)$.
We compare the performance of the proposed CLO scheme with resource allocation strategies based on the conventional LO framework. Recall that LO addresses only long-term constraints characterized by averages or higher-order statistical moments. Two primary baseline strategies are considered: 1) standard LO strategies tailored to long-term average constraints [8], and 2) LO strategies designed to handle outage probability constraints [54]. In the standard LO, we replace the deterministic constraint (23) by the average constraint [8]
@@ V-D Comparison with LO-based Resource Allocation Strategies
For the second benchmark, we impose the following outage probability constraint
$\lim_{F\to\infty}\frac{1}{F}\sum_{f=0}^{F-1}\mathbb{P}_{r}(\overline{L^{k}_{f}}>L^{k}_{\max})\leq\epsilon^{k},$ (38)
where $L^{k}_{\max}$ denotes the maximum tolerable reliability per frame, and $\epsilon^{k}$ specifies the target long-term outage probability. LO guarantees long-term reliability constraints by reformulating them as queue stability conditions associated with virtual queues for (37), and (38) [8]. We refer to Section II of the supplementary materials for further details.
Accordingly, while CLO updates the reliability hyperparameters $\theta^{k}$ at the end of each time frame (every $S$ time slots), the competitive LO formulations treat them as variables to be optimized at each time slot $t$. Treating these variables as discrete within the set $\{0.1,0.2,\dots,0.9\}$ yields a mixed-integer optimization problem, whose complexity grows exponentially with the number $K$ of users. Thus, to make the LO problems computationally feasible, in each slot we force all the users to employ the same threshold, i.e., $\theta^{k}(t)=\theta^{*}(t)$.
@@
The reliability constraints are highlighted by a dotted blue line in Figures 8(b) and 8(d), which plot the FNR evolution versus time for 30 tasks realizations. These plots are obtained under a comparable energy consumption for the two optimization strategies, which corresponds to the rightmost points in Figure 8(a). The continuous blue curves in Figures 8(b) and 8(d) highlight the worst-case FNR upper bound, computed by setting $M=1$ in (33), while the red curves identify an (a posteriori) upper bound, obtained by estimating the value of the constant $M$ among 30 realizations of CLO.
@@
Figure 8(c) shows two FNR outage curves for LO, defined as the probabilities to violate during convergence the worst case and the a posteriori upper bound, of CLO. The curves are obtained by evaluating the fraction of realizations (among 50), with an FNR value above the upper bound in Proposition 1, deterministically guaranteed by CLO for any realization. While CLO consistently remains within the theoretical bounds, LO exhibits a high likelihood to exceed them, with a probability that increases over time as the bound gets tighter. Comparisons with LO with Outage Probability Control: To ensure a fair comparison with LO strategies that incorporate outage probability control, we set the threshold value as $L^{k}_{\max}=r^{k}+r^{k}/10$ in (38), corresponding to a 10% margin above the target reliability $r^{k}=0.13$, for all users. With this choice, we evaluated the empirical outage probability achieved by CLO, defined as the frequency with which $\overline{L^{k}_{f}}$ exceeds $L^{k}_{\max}$, averaged over 30 independent realizations. The resulting average was $\epsilon^{k}\approx 32\%$, which was then adopted as the outage probability target in the outage constraint (38).
Figure 8(a) illustrates that the LO strategy with outage probability constraints (yellow curve) yields the worst energy–precision trade-off among the compared methods. This result stems from the fact that, in order to satisfy the outage probability constraint, LO tends to prioritize reliability over precision.
This observation is corroborated by Figure 9, which presents the long-term reliability achieved by the three competing strategies. The plot shows the average long-term reliability loss, with results averaged over 30 independent realizations of the task sequence. Unlike the standard LO and the proposed CLO scheme, the LO strategy with outage probability control consistently exhibits lower long-term reliability loss. This, due to the intrinsic trade-off between precision and reliability, also leads to diminished precision performance. Furthermore, it is observed that the CLO scheme achieves the lowest standard deviation among all the three strategies, highlighting its advantage in attaining a more stable solution in terms of long-term reliability.
V-E Effect of the Trade-off Parameter $V$
(a) Average energy consumption and precision loss vs. parameter $V$ ($d^{k}=0$). See Figure 10(b) for the color legend.
(b) Average energy vs. latency trade-off.
Figure 10: Behavior of CLO for different values of the Lyapunov trade-off parameter $V$: (a) Average energy consumption and precision loss as a function of $V$; (b) Energy vs latency trade-off.
In this section, we investigate the impact of the Lyapunov trade-off parameter $V$ (see (31)) on the network cost under both average latency and strict reliability constraints. As established in Lyapunov optimization theory, the parameter $V$ plays a crucial role in balancing performance and queue stability. Specifically, as the parameter $V$ increases, the average cost achieved by LO deviates from the optimum by an additive error of order $\mathcal{O}(1/V)$, while the average queue size grows proportionally to $\mathcal{O}(V)$ [8].
To highlight the effect of the trade-off parameter $V$ on CLO, we consider the single-hop edge-inference scenario depicted in Figure 5(a) with edge-to-edge latency constraints. Following the methodology proposed in [29], we augment problem (II-H) by incorporating an average constraint on the total queue length for each user, defined as
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{Q^{k}_{\mathrm{tot}}(t)\}\leq Q^{k}_{\mathrm{avg}}\hskip 20.0pt\forall k,$ (39)
where $Q^{k}_{\mathrm{tot}}(t)=Q^{k}_{k}(t)+Q_{s}^{k}(t)$, and $Q_{s}^{k}(t)$ denotes the queue at the centralized edge server (i.e., node S4 in Figure 5(a)). This constraint can be interpreted as an average latency constraint. Indeed, assuming a constant task arrival rate $\overline{A^{k}}=\lambda^{k}/\delta$ (in tasks/sec) and a stationary queuing system, exploiting Little’s Law, the total latency and the average latency constraint can be written as $D^{k}_{\mathrm{tot}}(t)=Q^{k}_{\mathrm{tot}}(t)/\overline{A^{k}}$, and $D^{k}_{\mathrm{avg}}=Q^{k}_{\mathrm{avg}}/\overline{A^{k}}$ respectively.
We consider a simulation time $T=10,000$ slots of $\delta=10$ ms and a reliability loss constraint $r^{k}=0.15$. Reliability is computed over frames composed of $S=10$ slots. The inference tasks are encoded with 8 bit per pixel, resulting in a task size $W^{k}=192$ KB, and generated according to an i.i.d. Bernoulli distribution with probability $\lambda^{k}=0.8$ for all the users. We impose a queue length constraint $Q^{k}_{\mathrm{avg}}=4$ tasks/slot, equivalent to a latency constraint $D^{k}_{\mathrm{avg}}=50$ ms.
To investigate the impact of the trade-off parameter $V$ on the precision/reliability balance, we first rewrite the cost function in (26) as
$(1-\beta)E_{\mathrm{tot}}(t)+\beta F_{\mathrm{tot}}(t),\quad\beta=\frac{\eta}{1+\eta},$ (40)
where the parameter $\beta\in[0,1]$ regulates the trade-off between energy consumption and precision loss. We evaluate the strategy for $\beta\in\{0.1,0.5,1\}$ and $V\in\{1,5,10,50,100,\dots,1000\}$.
Figure 10(a) reports the average energy consumption and precision loss as functions of the trade-off parameter $V$. Each curve corresponds to a fixed value of the weighting parameter $\beta$, with $V$ varying across the specified range. Results are obtained by averaging over the last 1,000 time slots, after convergence. As $V$ increases, the average precision loss decreases, with a consequent higher energetic consumption. On the other hand, higher values of $\beta$ result in improved precision due to more frequent task offloading to the edge server, which also leads to higher energy consumption.
Figure 10(b) illustrates the trade-off between energy consumption and latency. It can be observed that the optimization strategy consistently satisfies the average latency constraint, which is indicated by the red dashed line. Specifically, as the trade-off parameter $V$ increases, both energy consumption and latency increase, until the long-term latency constraint is tightly met. This behavior confirms that larger values of $V$ lead to a higher congestion state in the system.